OpenStudy (idealist10):

How to simplify sqrt(4-y^2)*y-sqrt(2y-y^2)*y?

OpenStudy (idealist10):

\[\sqrt{4-y^2}y-\sqrt{2y-y^2}y\]

OpenStudy (idealist10):

@ganeshie8 @satellite73 @thomaster @e.mccormick

OpenStudy (anonymous):

\[\begin{align*} y\sqrt{4-y^2}-y\sqrt{2y-y^2}&=y\sqrt{2-y}\sqrt{2+y}-y\sqrt{y}\sqrt{2-y}\\\\ &=y\sqrt{2-y}\left(\sqrt{2+y}-\sqrt y\right) \end{align*}\] I don't think there's much more you can do with this. It's not really simplifying, either...

OpenStudy (idealist10):

You're right, man. Thanks.